OpenStudy (anonymous):

1.Find the 13th partial sum of ( see drawing) a 118 b 708 c 1,534 d 767

OpenStudy (anonymous):



OpenStudy (anonymous):

ok this is \[\sum_{i=1}^{\infty}9i-4\] right?

OpenStudy (anonymous):

there is not a chance in hell this converges, because \(9>1\) and so this gets larger and larger. we can however find the partial sums

OpenStudy (anonymous):

you are asked for \[\sum_{i=1}^{13}9i-4\] right? that is the 13th partial sum

OpenStudy (anonymous):

I am not really sure, I just know I have to find the 13th partial sum of that sigma notation and i have no clue how to find partial um

OpenStudy (anonymous):

and there are a couple of different ways to compute it. someone else might have a snappy way, but i use this \[\sum_{i=1}^ni=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

ok lets go slow first of all is it clear what \[\sum_{i=11}^{13}9i-4\] means?

OpenStudy (anonymous):

What is \[\sum_{i=1}^ni=\frac{n(n+1)}{2}\] woudl you mind drawing it out for me?

OpenStudy (anonymous):

\[\sum_{i=1}^{13}9i-4\]means add, starting with \(i=1\) then \(i=2\) etc until you get up to \(i=13\)

OpenStudy (anonymous):

ohh, okay

OpenStudy (anonymous):

Satellite's correct, the 13th partial sum is just taking n from 1 to 13. So you're just summing the first 13 values of the sequence.